OpenStudy (anonymous):

4) A bullet of mass 12 grams is fired into a large block of mass 1.5 kg suspended from light vertical wires. The bullet imbeds in the block and the whole system rises 10 cm. Find the velocity of the bullet just before collision.

OpenStudy (anonymous):

m=.012 kilograms M=1.5 kilograms Height= 0.1meters find velocity BEFORE collision

OpenStudy (anonymous):

loss in kinetic energy = gain in potential energy 0.5mv^2 - 0 = mg(h+0.1) - mgh this should do it. but i'm not sure about it.

OpenStudy (anonymous):

how bout using the formula of conservation of momentum or conservation of energy?

OpenStudy (anonymous):

conservation of momentum will be hard to utilize cause we can't figure out the total momentum after collision. but we are provided here with the increase in potential energy after the collision so using conservation of energy is better .

OpenStudy (noelgreco):

You don't need to calculate the total momentum after the collision. You need 2 separate equations. Since momentum is conserved even in inelastic collisions such as this, mv = (m + M) v' where m is the mass of the bullet, M is the mass of the block, v is the velocity of the bullet, and v' is the velocity of the bullet and block an instant after the collision. That's the first equation.